I need help with an age problem. Mrs.Bruce lacks 1 year from being 5 times as old as her son. Five years from now she will lack 1 year from being 3 times as old as her son will be then. Find each of their ages.
Answer by mananth(16946)

son's age =x
Mrs. Bruce's age = 5x-1
5 years later
son = x+5
Mrs. Bruce = 5x-1+5=5x+4
5x+4=3(x+5)-1
5x+4=3x+15-1
5x+4=3x+14
5x-3x=10
2x=10
/2
x=5
son's age =5
Mother's age = 5x-1 = 25-1=24 years